After we took that noise spectrum we started putting in the PD for the transmission readout. It can only take 55uW according to the spec sheet so we started by turning the laser power down a bit and then we were putting in a BS to reduce the power even more. We were just getting this set up when the cavity stopped locking properly. I was working on the theory that it was the lower laser power was causing issues, so I turned it back up - still won't lock. I checked the beam is on the reflection PDs (it is) and that they weren't saturating (they're not).
We're getting big dips in power on the reflection PD when I scan the cavity through resonance. We also get a big error signal out of the mixer when I scan the cavity, but it looks kind of one sided (there is a massive positive side and hardly any negative side). I can also see the error signals from the sidebands resonating, and they look more normal.
What we're seeing when we try to lock is that it only stays locked for a couple of seconds, and we get big fluctuations in the control signal to the piezo, before it falls out of lock.
More investigation will ensue.
PS. The mojo beads were on the laser - Jenna has removed them in case they were giving it bad vibes. |